OpenStudy (anonymous):

Multiply (2n + 2)(2n - 2)

jimthompson5910 (jim_thompson5910):

Hint: use the difference of squares rule (a+b)(a-b) = a^2 - b^2

OpenStudy (anonymous):

I don't understand.

jimthompson5910 (jim_thompson5910):

ex: (7x + 5)(7x - 5) = (7x)^2 - (5)^2 (7x + 5)(7x - 5) = 49x^2 - 25

OpenStudy (dan815):